OpenStudy (anonymous):

If f(x)=e^x-(x^2)/2, then find the value of f(4)-f(0).

OpenStudy (anonymous):

simply input x=4 into the function and then x=0 in the function separately and minus the result of x=4 from the result of x=0

OpenStudy (anonymous):

given f(x)=e^x-(x^2/2 so now f(4)=e^4-4^2/2 =e^4-16/2 =>e^4-8 and now f(0)=e^0-0^2/2 =1-0 =1 so f(4)-f(0)=e^4-8-1 =e^4-9